# Towards General Conceptual Model Editing via Adversarial Representation Engineering

### v1

@@ Abstract
Recent research has introduced Representation Engineering (RepE) as a promising approach for understanding complex inner workings of large-scale models like Large Language Models (LLMs). However, finding practical and efficient methods to apply these representations for general and flexible model editing remains an open problem. Inspired by the Generative Adversarial Network (GAN) framework, we introduce a novel approach called Adversarial Representation Engineering (ARE). This method leverages RepE by using a representation sensor to guide the editing of LLMs, offering a unified and interpretable framework for conceptual model editing without degrading baseline performance. Our experiments on multiple conceptual editing confirm ARE’s effectiveness. Code and data are available at https://github.com/Zhang-Yihao/Adversarial-Representation-Engineering.
@@ 1 Introduction
Figure 1: An illustration of ARE framework. This example showcases how ARE can enhance the concept of "angry" within an LLM. The process involves an iterative dance between the generator and the discriminator. The generator produces outputs, while the discriminator refines its internal representation of "angry" based on these outputs. Through this back-and-forth training, the LLM gradually learns to produce outputs that align better with the concept of "angry."
@@
While Large Language Models (LLMs) have achieved remarkable success in a variety of tasks [37], their complex internal mechanism makes interpreting and censoring their behaviors (e.g., for safety alignment or hallucination reduction) challenging.
To improve the interpretability and consequently the safety of LLMs, numerous efforts have been dedicated to interpreting the internal mechanisms from various perspectives like feature attribution [11, 44], neuron analysis [16, 35], and self-explanation [20].
@@
Recently, Zou et al. proposed the idea of Representation Engineering (RepE[71], which offers a way of understanding how LLMs work internally by focusing on the overall feature representations rather than individual neurons. Specifically, RepE extracts and analyzes the intermediate features of different concepts (e.g., honesty, fairness, and harmlessness), enabling the monitoring of the internal behaviors of LLMs. More relevantly, RepE potentially allows editing and controlling the behaviors of LLMs by directly intervening in the internal hidden layers during inference.
@@
However, as RepE was essentially proposed to monitor the behaviors of LLMs, their proposed method for editing the model through representation vector incorporation is rather limited for practical uses. For instance, their method could disrupt the underlying structure of general LLMs, potentially hindering the model’s performance. Additionally, their representation vector used for model editing may not be accurate or robust, due to problems such as overfitting and adversarial samples. Furthermore, RepE might be heavily reliant on carefully chosen hyperparameters, and the naturalness and diversity of the LLM’s outputs could be negatively affected. To address these shortcomings, we investigate ways to efficiently fine-tune the model using the representations provided by RepE to achieve specific editing goals. Specifically, we attempt to train an oracle discriminator with the extracted representations given a particular goal of editing (e.g., harmlessness and trustfulness), then investigate how to use the discriminator to efficiently learn robust representations and subsequently edit the model accordingly. However, we found that the trained discriminator may (expectedly) fit non-robust features [21] and not be reliable for fine-tuning the models. Therefore, inspired by the adversarial learning networks [13], we extend our idea to conduct adversarial training between the generative model and the discriminator to improve the reliability of the oracle model.
@@
Motivated by these studies, we propose an Adversarial Representation Engineering (ARE) framework, utilizing the internal representations and adversarial learning from the generative model and the discriminator, as illustrated in Figure 1. ARE efficiently and effectively edits LLMs by leveraging representation engineering techniques. In each epoch, it performs two key steps. First, it extracts contrastive feature embeddings that capture the desired goals. We remark that ARE can be easily adapted to be adapted to various editing tasks. Secondly, it simultaneously trains both the LLM and the discriminator model. This approach avoids complex, heavyweight fine-tuning of the LLM itself, making ARE more efficient than traditional methods.
@@
We conduct extensive experiments to evaluate the effectiveness of ARE on various editing and censoring tasks, including editing the model to be better aligned or to remove existing alignment, or to always generate honest responses or dishonest ones. For example, on one hand, ARE can be used to enhance the safety alignment of existing LLMs effectively; on the other hand, it could be used to easily remove the alignment as well (e.g., for red-teaming). Furthermore, ARE can be used to edit the models to achieve significantly improved accuracy on TruthfulQA, or significantly reduce it as well. Additionally, we fine-tuned the model to achieve state-of-the-art TruthfulQA [28] accuracy. These results present strong evidence of the practicalities of ARE in terms of editing and censoring LLMs.
@@ 2 Related Work
Representation Engineering. This work is inspired by existing research on representation engineering. Numerous efforts have been put into extracting and visualizing the internal representations [45, 1] and concepts [33, 61] from AI models to improve their transparency and interpretability. In particular, the significant capability of LLMs has sparked great research interest in understanding their internal mechanisms [66, 46]. Representation engineering (RepE) [71], which seeks understanding and controlling representations of high-level cognition in LLMs, finds that there exist low-rank representations that can steer and control specific model capacity. Similar observations are also made in some specific scenarios, e.g. harmfulness [53, 68] and trustfulness [4]. However, RepE did not provide a practical solution to edit the model for practical goals.
Adversarial Learning. This work adopts adversarial learning to improve the robustness of representation learning. In response to the safety threats from adversarial examples [49, 13, 8], Adversarial training methods [32, 6, 42, 54] was first designed for defending deep neural networks (DNNs) against such adversarial attacks. By optimizing the min-max optimization objective with worst-case performance against adversarial examples, adversarial training successfully improves the robustness of DNNs and has proven to be one of the most effective technique [7, 3, 52] among various adversarial defense methods [15, 36, 63, 9]. In addition to the adversarial scenario, adversarial training also has the benefit of making the representation and prediction more reliable [14, 2] and interpretable [41, 47], thus also motivated other learning paradigms like domain generalization [12, 48] and contrastive learning [22, 67] for more robust representation. Our proposed adversarial representation engineering leverages the adversarial learning paradigm to make the oracle representation discriminator more robust and reliable.
Parameter-Efficient Fine-tuning. This work is related to parameter-efficient fine-tuning. Given the extremely large number of parameters in LLMs, parameter-efficient fine-tuning methods (PEFTs) are designed for tuning the LLM to be adapted to specific tasks with admissible computational overheads. Existing PEFTs can be mainly categorized as 1) module-based, which trains an extra small module in the model, like low-rank adaption (LoRA) [18, 30] and Adapters [17, 38]; and 2) prompt-based which optimizes a prompt or embedding for the task [43, 25]. While most PEFTs are designed for specific tasks, how to efficiently edit the model knowledge and style remains underexplored [34].
@@ 3 Notations and Problem Formulation
In this work, we focus primarily on LLMs, specifically decoder-only architectures, unless stated otherwise. These models are denoted as $M(\theta)$, where $\theta$ represents model parameters. The model $M$ is structured into several layers, collectively represented by the set $L\subseteq\mathbb{N}$, where each element $l\in L$ corresponds to the $l$-th decoder layer.
The model $M$ processes an input $x$, which is a prompt, to generate outputs as hidden states. Let $S$ denote the set of all valid input/output sentences. These hidden states are denoted by $H_{x}(\cdot)$, where $H_{x}(l)$ specifically refers to the hidden state from the $l$-th layer when model processes input $x$. Notably, each hidden state $H(l)$ is a vector in the $n$-dimensional real space $\mathbb{R}^{n}$. This architecture forms the basis for our analysis and further discussions in the paper. Moreover, the decoder model can generate responses, and we define the response generator function as $M_{\theta}(\cdot):S\to S$.
Next, we define a concept $c$ semantically. A concept specifically applies to sentences, or more precisely, to responses generated by the model $M$. To operationalize this, we introduce a function $J_{c}(\cdot):S\to\{0,1\}$ for each concept $c$. This function determines whether an input $s\in S$ embodies the concept $c$, as judged by a human expert. For example, for the concept "angry", the judgment function evaluates if a response is expressed in an angry manner; it outputs 1 if the response is furious, and 0 otherwise. For every concept $c_{+}$, there is a corresponding negation $c_{-}$, where the judgment function for $c_{-}$ is defined as the negation of that for $c_{+}$:
$J_{c_{-}}(s)=1-J_{c_{+}}(s)\text{ for all }s\in S.$
We are now ready to define the task of conceptual model editing. Assuming that the input prompts follow some implicit distribution $D$ defined on the space $S$, the task of conceptual editing, aimed at enhancing the concept $c_{+}$, is to fine-tune the model such that the response $r$ satisfies $J_{c_{+}}(r)=1$ for most inputs. This task is formally defined as
$\arg\max_{\theta}\mathbb{E}_{x\sim D}~{}J_{c_{+}}(M_{\theta}(x)).$
Notably, reducing the target concept $c_{+}$ is equivalent to enhancing the concept $c_{-}$. However, practically speaking, it is infeasible to edit in this manner due to the inability to access the true distribution $D$ or to implement the judgment function $J_{c}$ accurately. Therefore, a practical approach is to use a curated set of prompts to approximate these abstract concepts. To emulate the distribution, we use a small, diverse collection of prompts designed to cover a wide range of scenarios while avoiding an excessive representation of the target concept $c_{+}$, or alternatively, representing the negative concept $c_{-}$. This set of prompts is referred to as antithetical inputs, denoted $I_{A}$. Accordingly, our training objective becomes
$\arg\max_{\theta}\sum_{x\in I_{A}}J_{c_{+}}(M_{\theta}(x)).$
@@
$\sum_{x\in I_{A}}J_{c_{+}}(M_{\theta}(x))/|I_{A}|<\sum_{x\in I_{T}}J_{c_{+}}(M_{\theta}(x))/|I_{T}|.$
For example, consider the target concept "angry" that we wish to attain. To construct the antithetical inputs, we would gather a set of neutral prompts. Subsequently, to obtain the target inputs, we append the postfix "respond in an angry manner." to each prompt. This modification aims to reliably trigger responses that exhibit "anger", thereby constituting an effective set of target inputs.
As we utilize the target input set $I_{T}$ to illustrate the target concepts, the practical objective of fine-tuning shifts towards aligning the responses generated from $I_{A}$ as closely as possible with those from $I_{T}$. However, achieving token-level similarity is complex and overly fine-grained. Therefore, we employ a high-level approach known as representation engineering [71], which involves manipulating the representations—outcomes of an embedding function that maps the internal neural activities of a decoder model into the $n$-dimensional real space $\mathbb{R}^{n}$.
According to Zou et al. [71], these representations adeptly capture and reflect high-level conceptual attributes of the model in a structured, top-down manner. For any given concept $c$, it is possible to embed this concept within the representation space $\mathbb{R}^{n}$ as a distinct feature set apart within this Euclidean space. As illustrated in Figure 3(a), embeddings of opposite concepts are often distinctly separated, facilitating straightforward classification using simple classifiers.
The process of extracting these representations involves selecting tensors from the hidden states produced by processing an input $x$ across specified layers $L_{e}\subset L$. This process can be formally described by the mapping function $\mathcal{R}:[H_{x}(l)]_{l\in L_{e}}\to\mathbb{R}^{n}$, which transforms input space $S$ to representation space $\mathbb{R}^{n}$. A practical method for implementing this is to concatenate the hidden states from some selected layers.
By using these high-level representations, specifically target representations $R_{T}=\{\mathcal{R}(x)|x\in I_{T}\}$ and antithetical representations $R_{A}=\{\mathcal{R}(x)|x\in I_{A}\}$, we redefine our optimization goal. Representation serves as a proxy for the concept’s embedded features, enabling the definition of a similarity function $\mathcal{L}_{M(\theta)}(\cdot,\cdot)$ that quantifies the differences between these two sets of representations. The training objective is therefore established as
$\arg\min_{\theta}\mathcal{L}_{M(\theta)}(R_{T},R_{A}).$
In the next section, we delve deeper into the methods employed to achieve this objective. In particular, we show that the "loss function" $\mathcal{L}$ effectively functions as a discriminator.
@@ 4 Proposed Method
Contrary to the approach suggested in RepE [71], which focuses on generating a target representation vector that is simplistic and potentially overfitted, we introduce a representation discriminator designed to discern among representations so that we are able to learn a robust representation which allows effective model editing. This discriminator, embodied by a neural network, implicitly signifies the representation. By alternately updating this discriminator and the original model, we facilitate the development of a more refined and robust representation discriminator. This process forms the core of Adversarial Representation Engineering (ARE).
@@ 4.1 Adversarial Representation Engineering
Inspired by Generative Adversarial Networks (GANs) [13], ARE employs a dual-model design. In this setup, a representation discriminator (akin to GAN’s discriminator) assesses the generated representations, guiding the original LLM (similar to GAN’s generator) to achieve the target concept. We show this duality in Figure 2.
@@
Input: A representation generator decoder model $G$ that gets a string and returns corresponding representation, representation discriminator $D(\delta)$ with parameter $\delta$, targeted inputs $I_{T}=\{I_{T}^{(1)},\cdots,I_{T}^{(n)}\}$, antithetical inputs $I_{A}=\{I_{A}^{(1)},\cdots,I_{A}^{(n)}\}$, layers to edit $L_{e}$, layers to gather representations from $L_{r}$, optimizing epochs $T$, target concept label $y_{\texttt{Target}}$, learning rate for discriminator $l_{D}$

### v3

# Adversarial Representation Engineering: A General Model Editing Framework for Large Language Models

@@ Abstract
Since the rapid development of Large Language Models (LLMs) has achieved remarkable success, understanding and rectifying their internal complex mechanisms has become an urgent issue. Recent research has attempted to interpret their behaviors through the lens of inner representation. However, developing practical and efficient methods for applying these representations for general and flexible model editing remains challenging. In this work, we explore how to leverage insights from representation engineering to guide the editing of LLMs by deploying a representation sensor as an editing oracle. We first identify the importance of a robust and reliable sensor during editing, then propose an Adversarial Representation Engineering (ARE) framework to provide a unified and interpretable approach for conceptual model editing without compromising baseline performance. Experiments on multiple tasks demonstrate the effectiveness of ARE in various model editing scenarios. Our code and data are available at https://github.com/Zhang-Yihao/Adversarial-Representation-Engineering.
@@ 1 Introduction
Figure 1: An illustration of our proposed ARE framework. This example showcases how ARE can enhance the concept of "angry" within an LLM. The process involves an iterative dance between the generator and the discriminator. The generator produces outputs, while the discriminator refines its internal representation of "angry" based on these outputs. Through this back-and-forth training, the LLM gradually learns to produce outputs that align better with the concept of "angry."
@@
While Large Language Models (LLMs) have achieved remarkable success in a variety of tasks [34], their complex internal mechanism makes interpreting and censoring their behaviors (e.g., for safety alignment or hallucination reduction) challenging. To improve the interpretability and consequently the safety of LLMs, numerous efforts have been dedicated to interpreting the internal mechanisms from various perspectives like feature attribution [8, 40], neuron analysis [12, 33], and self-explanation [17].
@@
Recently, Zou et al. proposed the idea of Representation Engineering (RepE[69], which offers a way of understanding how LLMs work internally by focusing on the overall feature representations rather than individual neurons. Specifically, RepE extracts and analyzes the intermediate features of different concepts (e.g., honesty, fairness, and harmlessness), enabling the monitoring of the internal behaviors of LLMs. More relevantly, RepE potentially allows editing and controlling the behaviors of LLMs by directly intervening in the internal hidden layers during inference. However, as RepE was essentially proposed to monitor the behaviors of LLMs, their proposed method for editing the model through representation vector incorporation is rather limited for practical uses. For instance, their method could disrupt the underlying structure of general LLMs, potentially hindering the model’s performance. Additionally, the representation vector used for model editing may not be robust and heavily reliant on carefully chosen hyper-parameters, due to problems such as overfitting.
@@
To address these shortcomings, in this work we investigate ways to efficiently fine-tune the model using the representations provided by RepE to achieve specific editing goals. Specifically, we attempt to train an oracle discriminator with the extracted representations given a particular goal of editing (e.g., harmlessness and trustfulness), then investigate how to use the discriminator to efficiently learn reliable representations and subsequently edit the model accordingly. However, we found that the trained discriminator may (expectedly) fit non-robust features [18] and not be reliable for fine-tuning the models. Therefore, inspired by adversarial learning paradigms like GANs [10], we extend our idea to conduct adversarial training between the generative model and the discriminator to improve the reliability of the oracle model.
@@
Motivated by these studies, we propose an Adversarial Representation Engineering (ARE) framework, utilizing the internal representations and adversarial learning from the generative model and the discriminator, as illustrated in Figure 1. ARE efficiently and effectively edits LLMs by leveraging representation engineering techniques. In each epoch, it performs two key steps. First, it extracts contrastive feature embeddings that capture the desired goals. Secondly, it simultaneously trains both the LLM and the discriminator model. More details are discussed in the subsequent sections.
@@
We conduct extensive experiments to evaluate the effectiveness of ARE on various editing and censoring tasks, including editing the alignment and honesty abilities. Specifically, on one hand, ARE can be used to enhance the safety alignment of existing LLMs effectively; on the other hand, it could also be used to easily remove the alignment for red-teaming goals as well. Compared with some existing fine-tuning-based methods [37, 56], ARE can substantially decrease the refusal rate on harmful prompts from 20% to less than 1% on Llama2 [46]. Additionally, our ARE fine-tuned model can achieve the state-of-the-art TruthfulQA [26] accuracy. These results present strong evidence of the practicalities of ARE in terms of editing and censoring LLMs.
Our contribution in this work can be summarized as follows:
  1. 1.
    We propose the Adversarial Representation Engineering (ARE) framework, a novel and efficient method for model editing in LLMs. This framework enables flexible and bidirectional editing, facilitating both the enhancement and removal of specific concepts without compromising the baseline performance of the model.
  2. 2.
    We evaluated the ARE framework by addressing critical trustworthiness concerns such as alignment in jailbreak scenarios and hallucination control. Our framework successfully demonstrated its ability to both enhance model alignment and reduce hallucinations, providing robust solutions for improving LLM safety.
  3. 3.
    Our method sheds light on the internal workings of LLMs during the editing process, offering a level of transparency compared to traditional fine-tuning, which operates as a black box. By encoding concepts with contrasting sequence pairs and monitoring a small discriminator model, we gain insight into how specific layers are edited to maximize target representations, offering a more explainable approach to model tuning.
@@ 2 Related Work
Representation Engineering. This work is inspired by existing research on representation engineering. Since the significant capability of LLMs has sparked great research interest in understanding their internal mechanisms [64, 41, 47], Representation engineering (RepE) [69], which seeks understanding and controlling representations of high-level cognition in LLMs, has revealed that there exist low-rank representations that can steer and control specific model capacity. Similar observations are also made in some specific scenarios, e.g. harmfulness [48, 66] and trustfulness [2]. However, RepE did not provide a general solution to edit the model in a practical manner. Recently, the ReFT [53] (Representation Fine-Tuning) method has been proposed as a technique for fine-tuning models based on representation engineering. While ReFT has shown promise as a fine-tuning technique in the realm of representation engineering, it is not specifically designed to excel in model editing tasks.
Adversarial Learning. Our proposed method adopts adversarial learning intuitions to improve the reliability of representation discriminators. Adversarial training methods [30, 58, 49, 60, 32], which optimizes the min-max optimization objective with worst-case performance, was first designed for improving the robustness of machine learning models against adversarial examples [44, 10, 5]. In addition to the adversarial scenario, adversarial training has the benefit of making the representation and prediction more reliable [11, 1] and interpretable [38, 42], thus also been leveraged in other learning paradigms like image generation (GAN) [10], domain generalization [9, 43] and contrastive learning [20, 65] for more robust representation. Our proposed framework also leverages the adversarial learning paradigm to make the oracle representation discriminator more robust and reliable.
Parameter-Efficient Fine-tuning. This work is related to parameter-efficient fine-tuning. Given the extremely large number of parameters in LLMs, parameter-efficient fine-tuning methods (PEFTs) are designed for tuning the LLM to be adapted to specific tasks with admissible computational overheads. Existing PEFTs can be mainly categorized as (1) module-based, which trains an extra small module in the model, like low-rank adaption (LoRA[14, 28] and Adapters [13, 35], and (2) prompt-based, which optimizes a prompt or embedding for the task [39, 23]. While most PEFTs are designed for specific tasks, how to efficiently edit the model knowledge and style remains underexplored [31, 57].
@@ 3 Notations and Problem Formulation
In this work, we focus primarily on LLMs, specifically decoder-only architectures, denoted as $M(\theta)$ where $\theta$ represents model parameters. The model $M$ is structured into several layers, collectively represented by the set $L\subseteq\mathbb{N}$, where each element $l\in L$ corresponds to the $l$-th layer of the decoder.
Representations. During the model $M$ processes an input (prompt) $x$ to generate outputs, it also provides representations in the hidden layers. These hidden states can be formulated as $H_{x}(\cdot)$, where $H_{x}(l)\in\mathbb{R}^{n}$ specifically refers to the representation from the $l$-th hidden layer when the model processes input $x$. This architecture forms the basis for our analysis and further discussions in our work. Moreover, the response generated by the decoder model can be denoted as $M_{\theta}(\cdot):S\to S$, where $S$ denotes the set of all valid sentences.
Concepts. Next, we define a concept $c$ as the editing goal in the following. A concept applies to the responses generated by the model $M$. Specifically, we introduce a judge function $J_{c}(\cdot):S\to\{0,1\}$ to determine whether an input $s\in S$ aligns with the concept $c$ (ideally judged by human oracle). For example, for the concept "angry", $J_{c}(x)$ outputs $1$ if a response $x$ is expressed in an angry manner, and 0 otherwise. In addition, for every concept $c_{+}$, there is a corresponding negation $c_{-}$ which the judgment function is defined as the negation of that for $c_{+}$, i.e. $J_{c_{-}}(s)=1-J_{c_{+}}(s)\text{ for all }s\in S.$
Conceptual model editing. We are now ready to define the task of conceptual model editing. Assuming that the input prompts follow some implicit distribution $D$ defined on the space $S$, the task of conceptual editing, aimed at enhancing the concept $c_{+}$, is to fine-tune the model such that the response $r$ satisfies $J_{c_{+}}(r)=1$ for most inputs. This task is formally defined as
$\arg\max_{\theta}\mathbb{E}_{x\sim D}~{}J_{c_{+}}(M_{\theta}(x)).$ (1)
In general, it is infeasible to edit these concepts directly due to the inability to access the true distribution $D$ or to implement the judgment function $J_{c}$ perfectly. Therefore, a practical approach is to use a curated set of prompts to approximate these abstract concepts. This set of prompts is referred to as anti-target inputs, denoted $I_{A}$. Accordingly, our training objective becomes
$\arg\max_{\theta}\ \mathbb{E}_{x\sim I_{A}}J_{c_{+}}(M_{\theta}(x)).$ (2)
@@
$\mathbb{E}_{x\sim I_{A}}J_{c_{+}}(M_{\theta}(x))<\mathbb{E}_{x\sim I_{T}}J_{c_{+}}(M_{\theta}(x)).$ (3)
For example, consider the target concept of "anger" that we wish to attain (as illustrated in Figure 1). To construct the anti-target inputs, we would gather a set of neutral prompts. Subsequently, to obtain the target inputs, we append the suffix "respond in an angry manner." to each prompt. This modification aims to reliably trigger responses that exhibit "anger", thereby constituting an effective set of target inputs.
Representation extraction from concepts. Since we have utilized the target input set $I_{T}$ to illustrate the target concepts, the practical objective of fine-tuning shifts towards aligning the responses generated from $I_{A}$ as closely as possible with those from $I_{T}$. However, achieving token-level similarity is complex and overly fine-grained. Therefore, we employ a high-level approach known as representation engineering (RepE[69], which involves manipulating the representations, i.e. outcomes of an embedding function that maps the internal neural activities of each layer into the representation space $\mathbb{R}^{n}$. For any given concept $c$, the concept can be separated as a distinct feature set apart within this representation space of $\mathbb{R}^{n}$, as examplified in Figure 3(a). The process of extracting these representations involves selecting tensors from the hidden states produced by processing an input $x$ across specified layers $L_{e}\subset L$. This process can be formally described by the mapping function $\mathcal{R}:[H_{x}(l)]_{l\in L_{e}}\to\mathbb{R}^{n}$, which transforms input space $S$ to representation space as a subset of $\mathbb{R}^{n}$. A practical method for implementing this is to concatenate the hidden states from some selected layers.
By using these high-level representations, specifically target representations $R_{T}=\{\mathcal{R}(x)|x\in I_{T}\}$ and anti-target representations $R_{A}=\{\mathcal{R}(x)|x\in I_{A}\}$, we redefine our optimization goal. Representation serves as a proxy for the concept’s embedded features, enabling the definition of a similarity function $\mathcal{L}_{M(\theta)}(\cdot,\cdot)$ that quantifies the differences between these two sets of representations. The training objective is therefore established as
$\arg\min_{\theta}\mathcal{L}_{M(\theta)}(R_{T},R_{A}).$ (4)
In the next section, we delve deeper into the methods employed to achieve this objective. In particular, we show that the loss function $\mathcal{L}$ effectively functions as a discriminator.
@@ 4 Proposed Method
As discussed, the approach suggested in RepE [69] that focuses on generating a target representation vector may be unreliable and overfitted. To bridge this gap, we propose to train a representation discriminator to learn robust representations in an adversarial learning manner. This discriminator, embodied by a neural network, implicitly signifies the representation through the target concept. By iteratively updating this discriminator and the original model, we can facilitate a more refined and robust representation discriminator, forming the core of Adversarial Representation Engineering (ARE) as detailed in the following.
@@ 4.1 Adversarial Representation Engineering
Inspired by adversarial learning paradigms like Generative Adversarial Networks (GANs) [10], ARE employs a dual-model design. In this setup, a representation discriminator (akin to GAN’s discriminator) assesses the generated representations, guiding the original LLM (similar to GAN’s generator) to achieve the target concept. We show this duality in Figure 2.
@@
Input: A representation generator decoder model $G$ that gets a string and returns corresponding representation, representation discriminator $D(\delta)$ with parameter $\delta$, targeted inputs $I_{T}=\{I_{T}^{(1)},\cdots,I_{T}^{(n)}\}$, anti-target inputs $I_{A}=\{I_{A}^{(1)},\cdots,I_{A}^{(n)}\}$, layers to edit $L_{e}$, layers to gather representations from $L_{r}$, optimizing epochs $T$, target concept label $y_{\texttt{Target}}$, learning rate for discriminator $l_{D}$